Suppose that there are only two goods produced in the economy, Call center services and banking services. Prices (P), quantities (Q) and the number of workers (W) occupied in the production of each good for year 1 and year 2 are given by the following table : Year 1 Year 2 Two products P1 Q1 W1 P2 Q2 W2 Call centre 10 100 50 12 100 50 Banking 10 200 50 12 230 60 a. What is nominal GDP in each year? 


1
Expert's answer
2021-12-01T10:18:11-0500

a. The nominal GDP in each year is:

Y1 = 10×100 + 12×100 = 2200.

Y2 = 10×200 + 12×230 = 4760.